# Designing and Evaluating Hint Generation Systems for Science Education

### v1

@@ Abstract.
Large language models are influencing the education landscape, with students relying on them in their learning process. Often implemented using general-purpose models, these systems are likely to give away the answers, which could hinder conceptual understanding and critical thinking. We study the role of automatic hint generation as a pedagogical strategy to promote active engagement with the learning content, while guiding learners toward the answers. Focusing on scientific topics at the secondary education level, we explore the potential of large language models to generate chains of hints that scaffold learners without revealing answers. We compare two distinct hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to learners’ progress. Through a quantitative study with 41 participants, we uncover different preferences among learners with respect to hinting strategies, and identify the limitations of automatic evaluation metrics to capture them. Our findings highlight key design considerations for future research on hint generation and intelligent tutoring systems that seek to develop learner-centered educational technologies.
Do, Not, Us, This, Code, Put, the, Correct, Terms, for, Your, Paper
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
@@ 1. Introduction
Large language models have reshaped how students interact with digital learning tools. These models are increasingly used for everyday study practices, from gaining conceptual understanding to supporting problem-solving and writing (Levine et al., 2025; Rahma and Fithriani, 2024; Syarifah and Fakhruddin, 2024). At the same time, online learning continues to expand at unprecedented scales. According to UNESCO’s Global Education Monitoring Report (2023) (Antoninis et al., 2023), the number of students enrolled in massive open online courses reached at least 220 million in 2021, reflecting both the demand for and the accessibility of scalable educational resources. While prior work in education research has established a strong correlation between the student-teacher ratio and learning outcomes (Koc and Celik, 2015), teacher recruitment and retention crisis remains a worldwide conundrum (Huang et al., 2022; Williams III et al., 2022; Castro, 2022). ”Even where general teacher supply and demand are in balance, many countries face shortages of specialist teachers and shortages in schools serving disadvantaged or isolated communities.(Schleicher, 2012). These constraints create a growing reliance on automated support to bridge the gap in access to personalized instruction.
Despite their rapid adoption, many educational technologies are deployed with limited evaluation before reaching learners, and most evidence of the impact of education technology comes from the richest countries. Antoninis et al. (2023) reports that in the United Kingdom, only 7% of education technology companies had conducted randomized controlled trials, while a survey of teachers and administrators in 17 states within the United States revealed that only 11% requested peer-reviewed evidence prior to adoption. This context underscores the need for research that not only develops new forms of automated support, but also studies their effectiveness, limitations, and design trade-offs.
@@
Recent advances in LLM reasoning have spurred exploration of their potential for education, continuing a trajectory that began well before LLMs with intelligent tutoring systems (ITS) (Chu et al., 2025; Alhafni et al., 2024; Yan et al., 2024; Abd-Alrazaq et al., 2023). However, as in the pre-LLM era, the majority of research has centered on structured domains such as programming and mathematics, where correctness can be precisely defined and evaluation is straightforward. In contrast, relatively little attention has been paid to unstructured problem spaces, such as science learning, where solutions may be expressed in varied ways. To address this gap, our work focuses on hint generation systems for scientific problem solving, specifically designing hints for factoid questions as the first step towards scaffolding learners to enhance understanding and critical thinking in science domains (see Figure 1).
Figure 1. Illustration of automatic chain-of-hint generation for scientific question answering.
@@
We introduce the concept of a chain-of-hints, where information is revealed incrementally to scaffold the learner’s reasoning process. Inspired by the principles of human tutoring, this approach mirrors how teachers provide just-in-time interventions — offering minimal guidance first, and providing progressively more explicit support as needed (Wood et al., 1976). We systematically compare two strategies for generating and delivering these hints: pre-generated static hints and adaptive dynamic hints that respond to learner’s progress. Through this lens, we investigate how learners perceive the helpfulness of different hinting strategies, and derive design principles that can guide the development of pedagogically effective educational systems. We ground our study in the following research questions:
  • RQ1: What characteristics make hints useful to learners?
  • RQ2: How do learners perceive adaptive (dynamic) versus pre-generated (static) hinting strategies?
  • RQ3: To what extent do automatic evaluation metrics align with learners’ perceptions of hint quality?
In order to answer these questions, we conducted an evaluation study with 41 participants. The findings revealed the promise and complexity in designing effective hint generation systems. Learners valued hints that clarified concepts, provided additional context and incremental support. Yet, their preferences diverged over certain characteristics of static and dynamic hints, especially hints that took a context-agnostic approach to guide learners towards the answers. There is thus an implicit trade-off between short-term satisfaction and long-term accumulation of cognitive debt. Finally, we find that the automatic evaluation metrics are only partially aligned with the learner perceptions of hints’ helpfulness.
@@
Building on these insights, we discuss design implications for future hint generation and tutoring systems, applicable beyond factoid question answering. We call for learner-centered design practices that identify and account for diverse needs and preferences, rather than using general-purpose models. We also advocate for tutoring frameworks that integrate self-regulated learning with gradually faded support, enabling learners to develop metacognitive awareness and inhibiting over-reliance on automated assistance. In doing so, we position hint generation as a pathway to tackle the rising problem of cognitive debt due to over-reliance on readily available AI assistants designed to provide complete responses. This work makes three key contributions through our exploration of scientific hints as a form of interactive learning support: (1) we present results from an evaluation study with 41 learners toward understanding the components that make hints useful; (2) a suite of metrics for automatic evaluation of chain-of-hints across five dimensions; and (3) design implications for the development and evaluation of intelligent tutoring systems.
@@ 2.1. Automatic Hint Generation: Past and Present
Data-driven models offered a more scalable alternative, particularly in domains with well-structured problems such as programming (Barnes and Stamper, 2008; Rivers and Koedinger, 2017; Obermüller et al., 2021; Jin et al., 2012; Zimmerman and Rupakheti, 2015; Paaßen et al., 2018; Price et al., 2016; Rolim et al., 2017). These models typically followed a deterministic framework consisting of three components: (1) a corpus of diverse candidate solutions, (2) a matching algorithm to align learner attempts with the closest example, and (3) a path traversal algorithm to guide learners toward the correct solution. A rich body of surveys documents the innovations in this space (Le et al., 2013; Crow et al., 2018; McBroom et al., 2021; Mahdaoui et al., 2022). While effective in programming education, data-driven models were constrained by their dependence on high-quality training data, their limited adaptability beyond formal logic domains, and their difficulty in handling solution diversity.
@@
To overcome these limitations, recent work has increasingly explored LLM-based hint generation models. Large language models could offer various advantages, including improved reasoning ability, adaptability, and domain generalizability (Bommasani, 2021; Naveed et al., 2025; Kasneci et al., 2023; Minaee et al., 2024). While programming remains a popular testbed for LLM-based systems (Rogers et al., 2025; Demirtaş et al., 2025; Barros et al., 2025; Borchers and Shou, 2025; Ma et al., 2024; Liang et al., 2024; Kazemitabaar et al., 2024), the scope has expanded to other subjects with structured solution space, like mathematics. For instance, Wang et al. (2023b) curated a benchmark (ReMath) co-developed with teachers to evaluate step-by-step math tutoring, with the aim to use LLMs to assist math tutors in remediating student mistakes. Pal Chowdhury et al. (2024) proposed MWPTutor, a framework for generating math word problem hints by reformulating questions around the learner’s next step. Pardos et al. (2023) introduced OATutor, an open-source ITS platform to support adaptive tutoring research across multiple subjects. Building on this, Pardos and Bhandari (2024) evaluated ChatGPT-generated hints within OATutor, comparing them to human tutor hints across algebra and statistics. Their results demonstrate statistically significant learning gains from ChatGPT-generated hints.
@@
Recent work has also examined automatic feedback in more open-ended contexts. For instance, Lieb and Goel (2024) developed NewtBot, a LLM-based tutoring chatbot to support secondary school students in physics in German. They explored several pedagogy-grounded and context-grounded prompting strategies against a general purpose model, however, their findings reveal no statistically significant difference across the prompting strategies, likely due to small group size. In design education, Duan et al. (2024) utilized LLMs to generate heuristic feedback on UI mockups, offering constructive but sometimes inconsistent suggestions. In legal education, Weber et al. (2024) developed LegalWriter, an intelligent writing support system that can help novice students produce more structured and persuasive case solutions. These efforts highlight that while early work concentrated on structured topics, researchers are beginning to explore how LLM-based feedback can support learning in more open-ended domains.
@@
Despite this growing diversity, we identify two key reasons why research on hint generation has focused most heavily on structured domains. First, evaluation is more straightforward: solutions adhere to strict rules, allowing researchers to measure correctness with high reliability. Second, these domains align with the broader aim to improve formal reasoning capabilities of LLMs, developing benchmarks for establishing state-of-the-art performance (Chen et al., 2021b; Zhuo et al., 2024; Jiang et al., 2024; Paul et al., 2024; Glazer et al., 2024; Mishra et al., 2022; Li et al., 2020; Lu et al., 2023). By contrast, subjects involving a natural language solution space are underexplored within the intelligent tutoring systems community, due in part to the difficulty of evaluation. Through this work, we aim to contribute toward the development of hint generation methods that extend beyond formal logic settings.
@@ 2.2. Automatic Evaluation of Hints
In our work, we extended this line of research by investigating a broader set of pedagogically-grounded evaluation metrics for a chain of hints that better reflect the sequential nature of hint delivery. We also evaluated the alignment of these metrics to learners’ needs and experiences, identifying the blind spots of automatic evaluation. This analysis surfaces gaps in current metrics, and highlights open challenges for designing metrics that better capture hint utility from a learner-centered perspective.
@@
2.3. Learner-centered evaluation of generated hints
Third-party annotation. To assess the pedagogical quality of LLM-generated responses, Tack and Piech (2022) conducted a pairwise human-in-the-loop comparison study, with annotators rating responses along dimensions such as talking like teachers, understanding students, and helping improve student understanding. Their findings revealed that while LLMs often match teachers in conversational uptake, they lag significantly in their ability to foster deeper understanding.
In our work, we adopted a within-subjects randomized controlled trial to evaluate hints for scientific question answering. Our study emphasized on hint’s usefulness, insightfulness, and oversimplification, while also examining how these perceptions align or diverge from automatic evaluation metrics.
3. Scientific Hint Generation
Expert tutors are able to provide high quality support to students to help them meet their learning goals. These tutors take into consideration each individual’s learning objectives (Hoque, 2016; Sönmez, 2017) and prior knowledge (Orsmond and Merry, 2011; Biggs, 1996) to provide meticulously crafted direct instructions (Kozloff et al., 1999; Kim and Axelrod, 2005; Rosenshine, 2008) to guide the learners to the correct answer, molding them into deep conceptual thinkers (Rillero, 2016) with an enhanced sense of accomplishment (VanLehn, 2011; Merrill et al., 1992; Arroyo et al., 2014). Through this work, we aim to explore the capabilities of state-of-the-art reasoning models to provide this scaffolded guidance to learners to reach the correct answer. Limiting the focus to scientific hint generation, in this section, we formally define the chain-of-hints generation task (Section 3.1), and describe the dataset we use for the evaluation (Section 3.2) of several LLM-based baselines (Section 3.3). We also briefly describe the automatic evaluation metrics (Section 3.4) we use to select the model for user study, and provide the detailed description and analysis in Appendix A.
3.1. Problem Definition
[…]
$\mathbf{H}:\{q,a,\langle h_{1},h_{2}\ldots h_{i-1}\rangle\}\mapsto h_{i}$
where the next hint $h_{i}$ only depends on the question, the correct answer and the sequence of previously provided hints. In the dynamic setting, the function additionally conditions on the learner’s past incorrect attempts, i.e.,
Figure 2. Performance of best models from five baseline model families evaluated using automatic evaluation metrics. Orange (solid) bars denote the static hint generation results and blue (striped) bars denote the dynamic hint generation results.

### v2

# Designing and Evaluating Chain-of-Hints for Scientific Question Answering

@@ Abstract.
LLMs are reshaping education, with students increasingly relying on them for learning. Implemented using general-purpose models, these systems are likely to give away the answers, potentially undermining conceptual understanding and critical thinking. Prior work shows that hints can effectively promote cognitive engagement. Building on this insight, we evaluate 18 open-source LLMs on chain-of-hints generation that scaffold users toward the correct answer. We compare two distinct hinting strategies: static hints, pre-generated for each problem, and dynamic hints, adapted to a learners’ progress. We evaluate these systems on five pedagogically grounded automatic metrics for hint quality. Using the best performing LLM as the backbone of a quantitative study with 41 participants, we uncover distinct user preferences across hinting strategies, and identify the limitations of automatic evaluation metrics to capture them. Our findings highlight key design considerations for future research on tutoring systems and contribute toward the development of more learner-centered educational technologies.
Automatic Hint Generation, AI for Education, Evaluation Metrics
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Computer-assisted instructionccs: Computing methodologies Natural language generationccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Large language models (LLMs) have reshaped how students interact with digital learning tools. These models are increasingly used for everyday study practices, from gaining conceptual understanding to supporting problem-solving and writing (Levine et al., 2025; Sublime and Renna, 2024; Rahma and Fithriani, 2024; Syarifah and Fakhruddin, 2024). At the same time, online learning continues to expand at unprecedented scales, with at least 220 million students being enrolled in massive open online courses in 2021 (Antoninis et al., 2023). While prior work in education research has established a strong correlation between the student-teacher ratio and learning outcomes (Koc and Celik, 2015), teacher recruitment and retention crisis remains a worldwide conundrum (Huang et al., 2022; Williams III et al., 2022; Castro, 2022). These constraints create a growing reliance on automated support to bridge the gap in access to personalized instruction. However, despite their rapid adoption, many educational technologies are deployed with limited evaluation before reaching learners, and most evidence of the impact of education technology comes from the richest countries. According to UNESCO’s Global Education Monitoring Report (2023) (Antoninis et al., 2023), in the United Kingdom, only 7% of education technology companies had conducted randomized controlled trials, while a survey of teaching staff in 17 states within the United States revealed that only 11% requested peer-reviewed evidence prior to adoption.
Figure 1. Illustration of automatic chain-of-hint generation for scientific question answering (example interaction from the user study). After the learner’s initial incorrect response (“Worms”), the system provided a first hint directing attention to the kingdom (worms are not a biological kingdom) to which Candida and Trichophyton belong. When the learner then answered “Bacteria,” the system generated a second hint linking the organisms to familiar infections (e.g., athlete’s foot, thrush) without revealing the answer. Using this scaffolded guidance, the participant arrived at the correct response (“Fungus”), bridging the gap in the question and their prior knowledge.
@@
As LLMs become embedded in everyday study practices, there is a pressing need to evaluate not only what they generate for learners, but also how these outputs are assessed and integrated into instructional workflows (Chu et al., 2025; Alhafni et al., 2024; Yan et al., 2024; Abd-Alrazaq et al., 2023). However, the majority of research has centered on structured domains such as programming and mathematics, where correctness can be precisely defined and evaluation is straightforward (Jangra et al., 2024). In contrast, relatively little attention has been paid to unstructured problem spaces, with open-ended hint and solution space. To address this gap, our work focuses on hint generation systems for scientific question answering, specifically designing hints for factoid questions as the first step towards scaffolding learners to enhance understanding and critical thinking in science domains (see Figure 1). While factoid questions have well-defined answers that make evaluation tractable, the space of effective hints remains open-ended and diverse, requiring models to provide meaningful conceptual scaffolding. As such, factoid science questions offer a practical yet challenging testbed, bridging structured evaluation with realistic tutoring-style hint generation.
@@
In this paper, we introduce the concept of a chain-of-hints, where information is revealed incrementally (Section 3.1). Inspired by the principles of human tutoring, this approach mirrors how teachers provide just-in-time interventions — offering minimal guidance first, and providing progressively more explicit support as needed (Wood et al., 1976). We systematically compare two strategies for generating and delivering these hints: pre-generated static hints and adaptive dynamic hints that respond to learner’s progress. Through this lens, we investigate how learners perceive the helpfulness of different hinting strategies, and derive design principles that can guide the development of pedagogically effective educational systems. We ground our study in three research questions: RQ1: What characteristics of LLM-generated hint sequences do learners perceive as helpful or unhelpful during scientific question answering? RQ2: How do learners perceive adaptive (dynamic) versus pre-generated (static) chain-of-hints? RQ3: To what extent do automatic evaluation metrics align with learners’ perceptions of hint quality?
In order to answer these questions, we first conducted a comprehensive evaluation of 18 open-source LLMs (Section 3.3) using five pedagogically grounded automatic evaluation metrics: information gain, redundancy, consistency, readability, and leakage (Section 3.4). We observed substantial variance across model families and sizes, indicating that current LLMs are not uniformly capable of generating high-quality pedagogical feedback. We investigated the dependencies among metrics and observed notable trade-offs, including between information gain and leakage (Section 3.6). We selected the best-performing model at the Pareto-optimal front of the metric space for an exploratory study with 41 participants [^1] We intentionally recruited recent graduates rather than current high-school students to mitigate potential risks associated with exposing minors to experimental AI-based instructional systems (Yu et al., 2025). This study design choice may limit direct generalizability to in-school population (see Section 7). (Section 4). Quantitative results showed that access to hints enabled participants to answer approximately $\sim$22% more questions compared to a control baseline, with roughly $\sim$27% of hints rated as informative (Section 5.1). Qualitative feedback revealed that participants particularly valued hints that clarified underlying concepts, provided additional context, and offered incremental guidance (Section 5.2). However, their preferences diverged over certain characteristics of static and dynamic hints, static hints were valued for their contextual richness, whereas dynamic hints were useful for their answer-directed efficiency (Section 5.3). Additionally, we highlight critical limitations of the automatic evaluation metrics, that only partially aligned with the participants’ perceptions of hint helpfulness (Section 5.4).
@@
Building on these insights, we discuss design implications for future hint generation and tutoring systems, applicable beyond factoid question answering. We advocate for tutoring frameworks that integrate self-regulated learning with gradually faded support, enabling learners to develop metacognitive awareness and inhibiting over-reliance on automated assistance. In doing so, we position hint generation as a pathway to tackle the rising problem of cognitive debt due to over-reliance on readily available AI assistants designed to provide complete responses (Section 6).
This work makes three key contributions through our exploration of scientific hints as a form of interactive learning support: (1) a comprehensive evaluation of 18 open-source LLM using a five pedagogically-grounded evaluation metrics; (2) a detailed qualitative analysis from an exploratory study with 41 participants, aimed at identifying the components that make hints effective; and (3) design implications for the development and evaluation of intelligent tutoring systems, grounded in two distinct hint-generation strategies (static and dynamic).
@@ 2.1. Automatic Hint Generation: Past and Present
Data-driven models offered a more scalable alternative, particularly in domains with well-structured problems such as programming (Barnes and Stamper, 2008; Rivers and Koedinger, 2017; Obermüller et al., 2021; Jin et al., 2012; Zimmerman and Rupakheti, 2015; Paaßen et al., 2018; Price et al., 2016; Rolim et al., 2017). These models typically followed a deterministic framework consisting of three components: (1) a corpus of diverse candidate solutions, (2) a matching algorithm to align learner attempts with the closest example, and (3) a path traversal algorithm to guide learners toward the correct solution. A rich body of surveys have documented the innovations in this space (Le et al., 2013; Crow et al., 2018; McBroom et al., 2021; Mahdaoui et al., 2022). While effective in programming education, data-driven models were constrained by their dependence on high-quality training data, their limited adaptability beyond formal logic domains, and their difficulty in handling solution diversity.
@@
To overcome these limitations, recent work has increasingly explored LLM-based hint generation models. Large language models could offer various advantages, including improved reasoning ability, adaptability, and domain generalizability (Bommasani, 2021; Naveed et al., 2025; Kasneci et al., 2023; Minaee et al., 2024). While programming remains a popular testbed for LLM-based systems (Rogers et al., 2025; Demirtaş et al., 2025; Barros et al., 2025; Borchers and Shou, 2025; Ma et al., 2024; Liang et al., 2024; Kazemitabaar et al., 2024), the scope has expanded to other subjects with structured solution space, like mathematics. For instance, Wang et al. (2023a) curated a benchmark (ReMath) co-developed with teachers to evaluate step-by-step math tutoring, with the aim to use LLMs to assist math tutors in remediating student mistakes. Pal Chowdhury et al. (2024) proposed MWPTutor, a framework for generating math word problem hints by reformulating questions around the learner’s next step. Pardos et al. (2023) introduced OATutor, an open-source ITS platform to support adaptive tutoring research across multiple subjects. Building on this, Pardos and Bhandari (2024) evaluated ChatGPT-generated hints within OATutor, comparing them to human tutor hints across algebra and statistics using an in-context learning approach using chain-of-thought reasoning with self-consistency decoding (Wang et al., 2022). Their results demonstrate statistically significant learning gains from ChatGPT-generated hints.
While most prior work on LLM-generated hints relies on in-context learning–typically through few-shot examples or chain-of-thought prompting–recent research has begun exploring two complementary directions: synthetic data generation and post-training. In math tutoring, for example, Wang et al. (2025) demonstrate that LLMs can role-play teacher–student interactions to produce synthetic datasets, which can then be used to fine-tune smaller models (Zhang et al., 2025) or to train tutoring policies via LLM-based reward models (Dinucu-Jianu et al., 2025). These post-trained smaller models models can match or even exceed the performance of state-of-the-art systems on specialized tutoring tasks. Scarlatos et al. (2025) report that a Llama3-8B (Grattafiori et al., 2024) fine-tuned with direct preference optimization (Rafailov et al., 2023) outperformed GPT models on several mathematics-tutoring benchmarks. However, such gains come with substantial costs. Post-training requires significant computational resources and depends heavily on high-quality, pedagogically grounded datasets. As Dinucu-Jianu et al. (2025) highlight, even within the same problem-solving domain, different alignment strategies (e.g., supervised fine-tuning versus RL-based optimization) can lead to divergent outcomes.
@@
Recent work has also examined automatic feedback in more open-ended contexts. For instance, Lieb and Goel (2024) developed NewtBot, a LLM-based tutoring chatbot to support secondary school students in physics in German. They explored several pedagogy-grounded and context-grounded prompting strategies against a general purpose model, however, their findings reveal no statistically significant difference across the prompting strategies, likely due to small group size. Similarly, another small-scale investigation for conceptual physics problem solving by Wan and Chen (2024) found students to equally prefer GPT-generated feedback to human-written feedback. Expert human evaluators judged the LLM-generated responses to requiring minimal edits for most feedback. These efforts highlight that while early work concentrated on structured topics, researchers are beginning to explore how LLM-based feedback can support learning in more open-ended domains.
@@
Despite this growing diversity, we identify two key reasons why research on hint generation has focused most heavily on structured domains. First, evaluation is more straightforward: solutions adhere to strict rules, allowing researchers to measure correctness with high reliability. Second, these domains align with the broader aim to improve formal reasoning capabilities of LLMs, developing benchmarks for establishing state-of-the-art performance (Chen et al., 2021a; Zhuo et al., 2024; Jiang et al., 2024; Paul et al., 2024; Glazer et al., 2024; Mishra et al., 2022; Li et al., 2020; Lu et al., 2023). By contrast, subjects involving a natural language solution space are underexplored within the intelligent tutoring systems community, due in part to the difficulty of evaluation. Through this work, we aim to contribute toward the development of hint generation methods that extend beyond formal logic settings.
@@ 2.2. Automatic Evaluation of Hints
In our work, we extended this line of research by investigating a broader set of pedagogically-grounded evaluation metrics for a chain of hints that better reflect the sequential nature of hint delivery (Section 3.4). We also evaluated the alignment of these metrics to learners’ needs and experiences, identifying the blind spots of automatic evaluation (Section 5.4). This analysis surfaces gaps in current metrics, and highlights open challenges for designing metrics that better capture hint utility from a learner-centered perspective.
2.3. Learner-centered evaluation of generated hints
Third-party annotation. To assess the pedagogical quality of LLM-generated responses, Tack and Piech (2022) conducted a pairwise human-in-the-loop comparison study, with annotators rating responses along dimensions such as talking like teachers, understanding students, and helping improve student understanding. Their findings revealed that while LLMs often match teachers in conversational uptake, they lag significantly in their ability to foster deeper understanding. Similar findings were reported by Dai et al. (2023) for LLM-generated evaluations for an introductory data science course; where they found that while the LLM-generated outputs achieved higher readability scores compared to instructor-written responses, they did not translate into pedagogical accuracy. The LLM-generated assessments misaligned with the instructors’ evaluations of student work along key pedagogical dimensions of goal, benefit and clarity. For math error detection, expert tutor annotated analysis by Kakarla et al. (2024) revealed that while GPT-4 (Achiam et al., 2023) was able to accurately recognize the response criteria for student mistakes, it overidentified student errors for instances that human evaluators considered correct. While first-party feedback from learners can help gauge their affective state better, third-party evaluations using pedagogical experts can uncover aspects of learning that might be misrepresented by the learners’ feeling of learning (Deslauriers et al., 2019).
In our work, we adopted a within-subjects randomized controlled trial to evaluate hints for scientific question answering. Our study emphasized on hint’s satisfaction, informativeness, and answer leakage, while also examining how these perceptions align or diverge from automatic evaluation metrics. We describe the motivation for selecting these dimensions in our user study design (Section 4).
3. Scientific Hint Generation
We aim to explore the capabilities of state-of-the-art reasoning models to provide scaffolded guidance to users to reach the correct answer. Limiting the focus to scientific hint generation, in this section, we formally define the chain-of-hints generation task (Section 3.1), and describe the dataset we use for the evaluation (Section 3.2) of several LLM-based baselines (Section 3.3). We also briefly describe the automatic evaluation metrics (Section 3.4) we use to select the model for user study, and provide the detailed description and analysis in Appendix A.
3.1. Problem Definition
$P(a|q,h_{i})<1\;\forall i\in\{1,2,...,k\}$
$\mathbf{H}:\{q,a,\langle h_{1},h_{2}\ldots h_{i-1}\rangle\}\mapsto h_{i}$
where the next hint $h_{i}$ only depends on the question, the correct answer and the sequence of previously provided hints. In the dynamic setting, the function additionally conditions on the user’s past incorrect attempts, i.e.,